# Managing Classical Processing Requirements for Quantum Error Correction

### v1

@@ 7. Evaluations
Figure 9. (a) Max. concurrent critical decodes for all layouts and workloads; (b) Total execution time (in code cycles) needed to execute all workloads with different layouts (c) Estimated number of logical qubits for all workloads (code distances estimated by Azure QRE (Beverland2022, ) also annotated).
@@ 7.1. Research Questions
    How long do qubits go undecoded when using different scheduling policies?
@@
    How do scheduling policies affect memory usage when using virtualized decoders?
@@ 7.2. Baseline Statistics
[…]
@@ 7.3. Decoder Scheduling Efficacy
Figure 10. Hardware decoders used for different configurations (EDPC layout).
@@
Fig. 10 shows the three decoder configurations selected for this work.
  • Max. Concurrency refers to the configuration where the number of hardware decoders in the system corresponds to the peak concurrent critical decodes for every workload shown in Fig. 9(a).
The minimum concurrent critical decodes corresponds to two critical decodes between two logical qubits. All evaluations are for algorithmic logical qubits, logical qubits required for distillation are not considered. For the All Qubits configuration, the longest undecoded sequence length will be zero, since every logical qubit has an assigned decoder.
@@ 7.3.1. Longest Undecoded Sequences
To evaluate the performance of the decoder scheduling policies described in Section 4, we determine the longest undecoded sequence lengths for all workloads when using the Max. Concurrency and Midpoint configurations. Since these configurations use far fewer hardware decoders than qubits, the longest undecoded sequence is a good measure of whether qubits are being starved of decoding. Fig. 11(b) shows the longest undecoded sequence lengths for the (a) Max. Concurrency and (b) Midpoint configurations. The MFD policy leads to qubits being starved of decoding, since it prioritizes qubits that have frequent critical decodes. For the Max. Concurrency configuration, the gndstate and qaoa workloads do relatively well with the MFD configuration signifying that almost all logical qubits have a similar number of critical decodes, leading to a fairer scheduling. While the RR policy performs significantly better than the MFD policy, the MLS policy consistently performs better than both policies – MLS reduces the longest undecoded sequence lengths for almost every workload to $\sim 10$ slices for both Max. Concurrency and Midpoint configurations.
(a) Max. Concurrency
(b) Midpoint
Figure 12. Peak memory required for storing undecoded syndromes for different scheduling policies when using the (a) Max. Concurrency; (b) Midpoint configurations.
@@ 7.3.2. Memory Usage
The reduction in the longest undecoded sequence lengths also corresponds to lower memory usage for storing undecoded syndromes. This is crucial since reducing the number of hardware decoders will require more memory to store syndromes for qubits that have not been decoded. Fig. 12(b) shows the memory required for different workloads with the (a) Max. Concurrency and (b) Midpoint configurations (the Azure QRE estimated the code distances used to determine the memory requirements). Due to longer undecoded sequences, the MFD policy can require up to 100 GB of memory for some workloads while the MLS policy rarely requires more than 100 MB of memory, which is orders of magnitude better than the MFD policy and 2-4x better than the RR policy.
Figure 13. Total slices required by workloads with different decoder configurations normalized to the slices required by the baseline (MLS policy).
[…]
[some figure]
7.3.3. Slowdown due to Fewer Decoders
Fewer hardware decoders imply that some critical decodes in an arbitrary slice have to be deferred to subsequent slices, resulting in potentially more slices for completing the program. Fig. 13 shows the number of slices required for all workloads normalized with respect to the baseline number of slices shown in Fig. 9(b) for the Min. Concurrency, Max. Concurrency, and Midpoint configurations. Since the Min. Concurrency configuration allocates only four decoders for two critical decodes per slice, some workloads are slowed down by ¿10%. The Midpoint configuration however does not cause any slowdown except in the gndstate workload, and there is no slowdown for the Max. Concurrency configuration.
Figure 14. Increase in the LER (lower is better) with the Midpoint configuration.
[…]
[some figure]
7.3.4. Impact on Logical Error Rate
As discussed in Section 3, delaying decoding does not affect the logical error rate (LER) by itself. However, leaving a qubit undecoded for a long time before a critical decode can slow the decoder down, thus requiring additional rounds to complete the computation. In this evaluation, we would thus like to show how the longest undecoded sequence length can impact the LER. Fig. 14 shows how the MFD policy can increase the final LER of the computation due to longer undecoded sequence lengths. The RR policy increases the LER slightly for adder-28 and random-40, and the MLS policy does not incur any degradation in the LER. Note that this estimation is optimistic since we assumed the number of rounds increases linearly with the undecoded sequence length – in reality, it could be worse since the decoder latency can increase exponentially with the undecoded sequence length.
Figure 15. Reduction in the peak memory usage when using the Midpoint configuration with software offloading.
@@ 7.3.5. Software Offloading
The longest undecoded sequences (and consequently the memory requirements) can be reduced further by leveraging software decoders. The only constraint while doing so is that due to longer software decoding latencies, critical decodes should not be delayed because prior software decodes for a logical qubit have not yet finished. For evaluating the effect of software decoding, we set the number of hardware decoders to the Midpoint configuration and make a pessimistic assumption that a single slice worth of syndromes takes three slices (about 3$\times$×d microseconds) to be decoded in software (in reality, it could be much lower with optimized software decoders (Higott2023, )). Fig. 15 shows the reduction in the peak memory usage for all scheduling polices when software offloading is performed – software offloading can achieve a reduction of up to 3x.
@@ 7.4. Discussion
The results shown in previous sections show that VQD reduces the number of hardware decoders for algorithmic logical qubits by nearly one order of magnitude for most workloads with the Midpoint configuration, which, when combined with the MLS scheduling policy, results in significantly reduced memory requirements and low undecoded sequence lengths. Memory requirements and undecoded sequence lengths can be further reduced by offloading some non-critical decodes to software.
[…]
Applicability to other codes: While our evaluations have focused on the Surface Code, other error-correcting codes such as quantum LDPC codes (Bravyi2024, ) also require decoding with a significantly higher complexity than Surface Code. qLDPC codes use Belief Propagation for decoding errors (Old2023, ) along with Ordered Statistics Decoding (OSD) (Roffe2020, ; Fuentes2021, ) to enable accurate decoding. These algorithms are complex and highly resource intensive (Gong2024, ). Building fast, accurate decoders for such codes will likely require significant hardware resources. Our work enables amortizing the cost of expensive decoders via virtualization to build efficient and scalable quantum memory using qLDPC codes.
Better Capacity Planning: We envision large-scale quantum computers will be closely integrated with HPC-style systems, where scientific applications can leverage quantum subroutines using QPUs (humbleQPU, ). In this setting, non-critical software decoders can run on traditional HPC platforms to alleviate the pressure on hardware decoders. Moreover, the virtualization of decoders can help us harness shot-level parallelism – all quantum programs, even on FTQC, must be executed multiple times. We can concurrently run the copies of quantum programs on multiple QPUs. However, quantum resources increase linearly for running “$k$” copies concurrently. Our work, VQD, shows that with decoder virtualization, we can enable effective sharing of classical resources, dramatically reducing overall costs and improving resource utilization.
@@ 8. Related Work
This is one of the first works to perform a workload-oriented study of the classical processing requirements and system-level scheduling policies for error-corrected quantum computers. Prior to this work, Bombín et al. (Bombin2023, ) introduced modular decoding, which is the closest work that divides the global decoding task to sub-tasks without sacrificing decoder accuracy. However, this work, and other works such as parallelized window decoding (Skoric2023, ; Tan2022, ) always assume decoders for every logical qubit. This work shows that not all qubits require access to fast decoders at all times, thus allowing decoders to be virtualized. Other works that are broadly connected to this work are summarized below.
System-level Studies: Delfosse et al. (Delfosse2023, ) studied the speed vs. accuracy tradeoff for decoders used in FTQC. XQSim (Byun2022, ) is a full-system FTQC simulator. Stein et al. (Stein2023, ) proposed a heterogeneous architecture for FTQC, virtual logical qubits were proposed in (Baker2021, ), Lin et al. (Lin2023, ) explored modular architectures for error-correcting codes and scheduling for distillation factories was proposed in (Ding2018, ). (Kim2024, ) described a blueprint of a fault-tolerant quantum computer.
Decoder Designs: Neural network based decoders (googleRNN, ; ueno2022, ; Overwater2022, ; Meinerz2022, ; Gicev2023, ; Varsamopoulos2017, ), LUT-based decoders (Das2022lilliput, ; Tomita2014, ), decoders based on the union-find algorithm (Das2022afs, ; Riverlane2023, ), and optimized MWPM decoders (Vittal2023, ; Alavisamani2024, ) have been proposed. In general, neural network decoders are generally far slower and therefore not ideal for fast qubit technologies such as superconducting qubits. Other predecoders (Delfosse2020, ; Smith2023, ) and partial decoders (Caune2023, ) have also been proposed. Decoders based on superconducting logic (Ueno2021, ; Ravi2023, ) target cryogenic implementations.
@@ 9. Conclusions
Scaling quantum computers to enable Quantum Error Correction will require specialized hardware for decoding errors. Prior work has focused on reducing the hardware resources required to build decoders. In this work, we take a full-system view and show that with the right decoder scheduling policy, it is not necessary for an error-corrected quantum computer to provide every logical qubit with a dedicated hardware decoder. The MLS policy enables the reduction of hardware decoders by up to 10x while requiring $\sim$100 MB or less of memory for storing undecoded syndromes without increasing the program execution time or the target logical error rate. The efficacy of the MLS policy is enhanced with software offloading of some decoding tasks. We also propose a noise-adaptive scheduling mechanism that can prioritize the decoding of logical qubits that incur a temporary increase in the physical error rate.

### v2

@@ 7. Evaluations
(a) Max. concurrent critical decodes.
(b) Total execution time (in code cycles)
(c) Estimated number of logical qubits (code distances estimated by Azure QRE (Beverland et al., 2022b) also annotated).
Figure 11. Baseline statistics.
@@ 7.1. Research Questions
    How long do qubits go undecoded when using virtualization?
@@
    How does virtualization affect Magic State Distillation?
  4. (4)
    How do error bursts affect decoder resource requirements?
  5. (5)
    How do slow decoders for qLDPC codes affect the efficacy of virtualization?
@@ 7.2. Baseline Statistics
We consider the baseline to have decoders for every logical qubit in the system. Figure 11(a) shows the maximum number of critical decodes that occur during the execution of different workloads for all selected layouts. The Compact layout has a maximum of one critical decode per slice between two logical qubits. Figure 11(b) shows the total time required to finish all workloads444shor-15 and qpe-50 were stopped prematurely due to long runtimes (shor-15 after 100,000 time slices for all layouts, qpe-50 after 72,000 time slices for Fast, EDPC, and 100,000 time slices for Compact). – the EDPC and Fast layouts are significantly faster than Compact, and the EDPC layout has a very slight advantage over Fast for most workloads and also uses fewer qubits 11(c) (only algorithmic logical qubits). Due to its higher degree of parallelism (due to its better use of routing lanes), we only consider the EDPC layout for all further evaluations.
@@ 7.3. Decoder Scheduling Efficacy
Figure 12 shows the three decoder configurations used in this work.
  • Max. Concurrency is the configuration where the number of hardware decoders in the system corresponds to the peak concurrent critical decodes for every workload shown in Figure 11(a).
Figure 12. Hardware decoders normalized with the Midpoint configuration (EDPC layout). Decoders used by the Midpoint configuration are annotated below the benchmarks.
@@
The minimum concurrent critical decodes corresponds to one critical decode between two logical qubits because a multi-body measurement merges the two patches into a single decoding task. We first focus our evaluations on algorithmic logical qubits, and then incorporate distillation factories as well. For the All Qubits configuration, the longest undecoded sequence length will be zero, since every logical qubit has an assigned decoder.
@@ 7.3.1. Longest Undecoded Sequences
To evaluate the performance of the decoder scheduling policies described in Section 4, we determine the longest undecoded sequence lengths for all workloads when using the Max. Concurrency and Midpoint configurations. Since these configurations use far fewer hardware decoders than qubits, the longest undecoded sequence is a good measure of whether qubits are being starved of decoding. Figure 13(b) shows the longest undecoded sequence lengths for the (a) Max. Concurrency and (b) Midpoint configurations. The MFD policy leads to qubits being starved of decoding, since it prioritizes qubits that have frequent critical decodes. While the RR policy performs significantly better than the MFD policy, the MLS policy consistently performs better than both policies – MLS reduces the longest undecoded sequence lengths for almost every workload to $\sim$10 time slices for both Max. Concurrency and Midpoint configurations.
Figure 14. Memory required (normalized by the MLS policy) for undecoded syndromes with the Midpoint configuration.
@@ 7.3.2. Memory Usage
The reduction in the longest undecoded sequence lengths also corresponds to lower memory usage for storing undecoded syndromes. Figure 14 shows the memory required for different workloads with the Midpoint configuration (memory requirements were determined using the code distances estimated by Azure QRE). Due to longer undecoded sequences, the MFD policy can require 3–4 orders of magnitude more memory than the MLS policy, which requires around 0.01MB of memory. The RR policy can require 3-10$\times$ more memory than the MLS policy.
Figure 15. Reduction in the longest undecoded sequence length when using the Midpoint configuration with software offloading.
@@ 7.3.3. Software Offloading
The longest undecoded sequences can be reduced further by leveraging software decoders. The only constraint while doing so is that critical decodes should not be delayed because prior software decodes for a logical qubit have not yet finished. For evaluating the effect of software decoding, we set the number of hardware decoders to the Midpoint configuration and make a pessimistic assumption that a single slice worth of syndromes takes three time slices (about 3$\times$×d microseconds) to be decoded in software (in reality, it could be much lower with optimized software decoders (Higgott and Gidney, 2023)). Figure 15 shows the reduction in the longest undecoded sequence length for all scheduling polices when software offloading is performed – software offloading can achieve a reduction of more than 30%.
Figure 16. Total reduction in decoders required when considering 15-to-1 MSD factories with the Midpoint configuration, EDPC layout. The number of MSD factories was fixed at 24 for all workloads, and the number of factories estimated by Azure QRE are annotated below the benchmark names.
[…]
[some figure]
7.3.4. Incorporating Magic State Distillation
So far, our evaluations have considered only algorithmic logical qubits. Since Lattice Surgery Compiler takes the qubit layout as an input, we fixed the number and locations of magic state distillation (MSD) factories in the EDPC layout at 24 for all workloads. This number was chosen to maintain a uniform EDPC layout across all workloads, and it is also the median number of MSD factories estimated by Azure QRE (annotated below the benchmarks names in Figure 16). 15-to-1 distillation is mapped to a Compact layout (Litinski, 2019a). Using this layout and the QASM circuit for a 15-to-1 MSD factory, the MLS policy was used to virtualize decoders. Table 1 shows the reduction in the number of decoders and the longest period for which a qubit remains undecoded for a single distillation factory. A 15-to-1 distillation factory requires five logical qubits (Litinski, 2019a) and with two decoders each for $X/Z$ observables if all qubits are supplied their own decoders. This can be reduced by 60% to 4 decoders with virtualization.
Table 1. Virtualizing decoders using the MLS policy can reduce the decoders required for 15-to-1 distillation without a substantial backlog of undecoded slices.
All Qubits After Virtualization
Longest Undecoded
Sequence Length
15-to-1 Distillation 10 decoders 4 decoders 4 time slices
Assuming that decoders are independent for every MSD factory, the reduction in the number of decoders can be computed using data from Figure 12 and Table 1, as shown in Figure 16. The MLS policy can reduce the total number of decoders required by up to 4$\times$. MSD is a probabilistic process, and factories will not be in sync. at all times. Thus, treating decoders for all factories to be independent is a conservative setting – in reality, it is likely that decoders can be virtualized among multiple factories, which can yield further reductions in the number of decoders in the system.
7.3.5. Dynamic Scheduling
Periodic events such as heating effects, cosmic rays, can result in a temporary increase in the error rate for some qubits. Such error bursts can necessitate dynamic scheduling of decoding for affected qubits. For evaluating the overheads of such error bursts on the virtualized decoders, we configured a burst probability $P_{Burst}$. For a workload requiring $N$ slices, $P_{Burst}\times N$ slices are randomly selected from which an active logical qubit is selected with a probability of $P_{Burst}$ to be decoded. Figure 17 shows the increase in the decoders required for different $P_{Burst}$ – dynamic scheduling only increases the decoder resource requirements for a 10% probability of a burst error happening. This shows that virtualization is effective even with moderate error burst probabilities.
Figure 17. Normalized increase in the number of decoders required with the Midpoint configuration due to error bursts with probability $P_{Burst}$.
[…]
[some figure]
Figure 18. (a) Total decoding task required to process the longest undecoded sequence of syndromes for all workloads, with the normalized decoder latency for the Surface Code (SC) as 0.5 and 0.99 for the qLDPC code. The Midpoint configuration was used for both cases; (b) Relative increase in the LER for different workloads for the SC and qLDPC setups.
[…]
[some figure]
7.3.6. Slowdown due to Decoding Latencies
Do slow and complex decoders, such as the ones used for qLDPC codes, work well with virtualization? Processing undecoded syndromes requires an additional number of rounds that is dependent on the decoding processing rate. For the Surface Code (SC), decoders that process syndromes significantly faster than the syndrome generation rate have been proposed (Vittal et al., 2023; Barber et al., 2023; Ziad et al., 2024). However, unlike the Surface Code, quantum LDPC codes (Bravyi et al., 2024; Roffe et al., 2020; Gong et al., 2024; Hillmann et al., 2024; Wolanski and Barber, 2024) have considerably slower decoders with software latencies in the order of milliseconds. To put that in context, software decoders for the Surface Code have microsecond latencies (Higgott and Gidney, 2023). Assuming accelerated decoders for qLDPC codes are designed, it is reasonable to expect them to be slower than Surface Code decoders. To see the impact of such slow decoders on virtualization, we modified our framework to support heterogeneous systems consisting of qLDPC blocks and Surface Code patches for compute (Stein et al., 2024). To do this, we assumed every magic state consumed involved an additional decoder to decode the ancillary system required for computation with qLDPC codes (Cross et al., 2024).
[…]
The results above show that the benefits of virtualization are only apparent when the decoding latency is much smaller than the syndrome generation latency. For both qLDPC and Surface Codes, it is thus highly desirable to have short decoding latencies, since this can enable more efficient virtualization. Decoders thus need not be designed for resource efficiency – they can be built for speed and accuracy. Resource savings can be attained at a system-level via virtualization.
@@ 7.4. Discussion
The results shown in previous sections show that VQD reduces the number of hardware decoders for algorithmic logical qubits by nearly one order of magnitude for most workloads with the Midpoint configuration, which, when combined with the MLS scheduling policy, results in significantly small undecoded sequence lengths of $\sim$10 time slices. With the inclusion of distillation factories, the total reduction in the number of decoders is up to 4$\times$.
Faster or Resource Efficient Decoders?: The faster a decoder is in relation to the syndrome generation latency, the more effective is decoder virtualization. Decoders with smaller hardware footprints (Vittal et al., 2023; Barber et al., 2023) generally sacrifice decoding accuracy for speed and hardware-efficiency. This work indicates that hardware-efficiency can be achieved with effective decoder virtualization on a system-level rather than at a module-level by leveraging fast, accurate decoders that need not be hardware efficient by themselves. This lends more motivation to the design of neural network decoders (Bausch et al., 2023; Acharya et al., 2024), which can be resource intensive but offer code distance independent decoding latencies and can leverage existing hardware (CPUs/GPUs) rather than require custom FPGA/ASIC solutions. This insight can be applied to build fast and accurate decoders for other codes such as qLDPC, Color codes as well.
Better Capacity Planning: We envision large-scale quantum computers will be closely integrated with HPC-style systems, where scientific applications can leverage quantum subroutines using QPUs (Britt and Humble, 2017). In this setting, non-critical software decoders can run on traditional HPC platforms to alleviate the pressure on hardware decoders. Moreover, the virtualization of decoders can help us harness shot-level parallelism – all quantum programs, even on FTQC, must be executed multiple times. We can concurrently run the copies of quantum programs on multiple QPUs. However, quantum resources increase linearly for running “$k$” copies concurrently. Our work, VQD, shows that with decoder virtualization, we can enable effective sharing of classical resources, dramatically reducing overall costs and improving resource utilization.
@@ 8. Related Work
This is one of the first works to perform a workload-oriented study of the classical processing requirements and system-level scheduling policies for error-corrected quantum computers. Prior to this work, Bombín et al. (Bombín et al., 2023) introduced modular decoding, which is the closest work that divides the global decoding task to sub-tasks without sacrificing decoder accuracy. However, this work, and other works such as parallelized window decoding (Skoric et al., 2023; Tan et al., 2022) always assume decoders for every logical qubit. This work shows that not all qubits require access to fast decoders at all times, thus allowing decoders to be virtualized. Other works that are broadly connected to this work are summarized below.
System-level Studies: Delfosse et al. (Delfosse et al., 2023) studied the speed vs. accuracy tradeoff for decoders used in FTQC. XQSim (Byun et al., 2022) is a full-system FTQC simulator. Stein et al. (Stein et al., 2023) proposed a heterogeneous architecture for FTQC, virtual logical qubits were proposed in (Baker et al., 2021), Lin et al. (Lin et al., 2023) explored modular architectures for error-correcting codes and scheduling for distillation factories was proposed in (Ding et al., 2018). (Kim et al., 2024) described a blueprint of a fault-tolerant quantum computer.
Decoder Designs: Neural network based decoders (Bausch et al., 2023; Ueno et al., 2022; Overwater et al., 2022; Meinerz et al., 2022; Gicev et al., 2023; Varsamopoulos et al., 2017; Acharya et al., 2024), LUT-based decoders (Das et al., 2022a; Tomita and Svore, 2014), decoders based on the union-find algorithm (Das et al., 2022b; Barber et al., 2023; Ziad et al., 2024), and optimized MWPM decoders (Vittal et al., 2023; Alavisamani et al., 2024) have been proposed. In general, neural network decoders are generally far slower and therefore not ideal for fast qubit technologies such as superconducting qubits. Other predecoders (Delfosse, 2020; Smith et al., 2023) and partial decoders (Caune et al., 2023) have also been proposed. Decoders based on superconducting logic (Ueno et al., 2021; Ravi et al., 2023) target cryogenic implementations.
@@ 9. Conclusions
Scaling quantum computers to enable Quantum Error Correction will require specialized hardware for decoding errors. Prior work has focused on reducing the hardware resources required to build decoders. In this work, we take a full-system view and show that with the right decoder scheduling policy, it is not necessary for an error-corrected quantum computer to provide every logical qubit with a dedicated hardware decoder. The MLS policy enables the reduction of hardware decoders by up to 10x without increasing the program execution time or the target logical error rate. The efficacy of the MLS policy is enhanced with software offloading of some decoding tasks. We also propose a noise-adaptive scheduling mechanism that can prioritize the decoding of logical qubits that incur a temporary increase in the physical error rate. Via virtualization, decoder systems can be made more scalable, and this insight can allow algorithm and hardware designers to focus on making decoders faster and more accurate, no matter the hardware cost.